Grade 12General Physics

Two thin lenses of power +7D and -3D are in contact. What is the focal length of the combination ?

Raghav
12 Years agoGrade 12

1 Answer

Saurabh Koranglekar
6 Years ago
Dear student

P=1/f

1/F = 1/f1 + 1/f2

Regards